A test of spelling ability is given to a random sample of 10 students before and after they completed a spelling course. The mean score before the course was 119.60 and after the course the mean score was 130.80. The standard deviation of the difference was 16.061. How would you test the hypothesis that the course significantly improved spelling performance?
Answer by stanbon(75887)

A test of spelling ability is given to a random sample of 10 students before and after they completed a spelling course. The mean score before the course was 119.60 and after the course the mean score was 130.80. The standard deviation of the difference was 16.061. How would you test the hypothesis that the course significantly improved spelling performance?
-----------
Ho: u(aft)-u(b4) <= 0
Ha: u(aft)-u(b4) > 0 (claim)
------
x-bar(aft) - x-bar(b4) = (130.8-119.6) = 11.2
-----
z(11.2) = (11.2-0)/16.061 = 0.6973
-----
p-value = P(z > 0.6973) = normalcdf(0.6973,100) = 0.2428
=================
Cheers,
Stan H.
==============================